40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Aug 5 11:04:07 2017, gautam, Update, SUS, MC1 glitches return MC1_glitch_Aug42017.pngMC1_glitch_zoomed.png
    Reply  Wed Aug 9 15:15:47 2017, gautam, Update, SUS, MC1 glitches return MC1_glitch.png
       Reply  Thu Aug 10 21:01:43 2017, gautam, Update, SUS, MC1 glitches debugging MC_QUAD_10AUG2017.jpgMCR_10AUG2017.jpg
          Reply  Fri Aug 11 12:32:46 2017, gautam, Update, SUS, MC1 glitches debugging MC1_freeswinging.pdfMC1_inmatrix.pngMC1_sensors.png
             Reply  Fri Aug 11 17:36:47 2017, gautam, Update, SUS, MC1 <--> MC3 MC1_glitch_watchdog_shutdown.png
                Reply  Sat Aug 12 14:09:36 2017, gautam, Update, SUS, Glitches stay on MC1 MC1_glitch.pngMC_12hr_trend.pngMC1_glitches_intermittent.png
                   Reply  Sat Aug 12 22:35:22 2017, rana, Update, SUS, Glitches stay on MC1 
                      Reply  Sun Aug 13 01:35:09 2017, Koji, Update, SUS, Glitches stay on MC1 
                         Reply  Mon Aug 14 20:01:38 2017, gautam, Update, SUS, Glitches stay on MC1 
                Reply  Wed Aug 16 19:50:17 2017, gautam, Update, SUS, MC1 <--> MC3 switched back 
                   Reply  Thu Aug 17 11:17:49 2017, gautam, Update, SUS, MC1 <--> MC3 switched back MC1_misaligned.pngMC1_glitch.png
                      Reply  Thu Aug 17 17:33:01 2017, gautam, Update, SUS, MC1 <--> MC3 switched back 
Message ID: 13195     Entry time: Fri Aug 11 12:32:46 2017     In reply to: 13187     Reply to this: 13196
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 glitches debugging 

Attachment #1: Free swinging sensor spectra. I havent done any peak fitting but the locations of the resonances seem consistent with where we expect them to be.

The MC_REFL spot appears to not have shifted significantly (so slow bias voltages are probably not to blame). Now I have to look at trend data to see if there is any evidence of glitching.

I'm not sure I understand the input matrix though - the matrix elements would have me believe that the sensing of POS in UL is ~5x stronger than in UR and LL, but the peak heights don't back that up.

Attachment #3: Second trend over 5hours (since frame writing was re-enabled this morning). Note that MC1 is still free-swinging but there is no evidence of steps of ~30cts which have been observed some days ago. Also, from my observations yesterday, MC1 glitched multiple times over a few hours timescale. More data will have to be looked at, but as things stand, Hypothesis #3 below looks the best.

Quote:
 

Some possible scenarios (assuming the free swinging spectra look alright and the various resonances are where we expect them to be):

  1. With the watchdog shutdown, the PIT/YAW bias voltages still goto the coil (low-passed by 4 poles @1Hz). So if the glitching happens in this path, we should see it in both the shadow sensors and the DC spot positions on the WFS.
  2. If the glitching happens in the shadow sensor readout electronics/cabling, we should see it in the shadow sensor channels, but NOT in the DC spot positions on the WFS (as the watchdog is shutdown, so there should be no actuation to the coils based on OSEM signals).
  3. If we don't see any glitches in WFS spot positions or shadow sensors, then it is indicative of the problem being in the coil driver board / dewhitening board/anti-aliasing board.
  4. I am discounting the problem being in the Satellite box, as we have switched around the MC1 satellite box multiple times - the glitches remain on MC1 and don't follow a Satellite Box. Of course there is the possibility that the cabling from 1X5/1X6 to the Satellite box is bad.

 

Attachment 1: MC1_freeswinging.pdf  42 kB  | Hide | Hide all
MC1_freeswinging.pdf
Attachment 2: MC1_inmatrix.png  5 kB  Uploaded Fri Aug 11 13:49:44 2017  | Hide | Hide all
MC1_inmatrix.png
Attachment 3: MC1_sensors.png  64 kB  Uploaded Fri Aug 11 17:52:48 2017  | Hide | Hide all
MC1_sensors.png
ELOG V3.1.3-